# MicroAdam: Accurate Adaptive Optimization with Low Space Overhead and Provable Convergence

### v1

@@ 4.1 Gradient and Error Compression
9:     $\theta_{t+1}=\theta_{t}-\eta\frac{\hat{m_{t}}}{\sqrt{\hat{v}_{t}+\epsilon}}$diamond Update the model parameters
@@ 4.2 Convergence Guarantees for General Smooth Non-convex Functions
The loss function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth and lower bounded by some $f^{*}\in\mathbb{R}$:
$\left\|\nabla f(\theta)-\nabla f(\theta^{\prime})\right\|\leq L\left\|\theta-\theta^{\prime}\right\|,\quad\text{for any }\theta,\theta^{\prime}\in\mathbb{R}^{d}.$
@@
For all iterates $t\geq 1$, the stochastic gradient $g_{t}$ is unbiased and uniformly bounded by some constant $G\geq 0$:
$\mathbb{E}[g_{t}]=\nabla f(\theta_{t}),\quad\left\|g_{t}\right\|\leq G.$
@@
For all iterates $t\geq 1$, the variance of the stochastic gradient $g_{t}$ is uniformly bounded by some constant $\sigma^{2}\geq 0$:
$\mathbb{E}[\|g_{t}-\nabla f(\theta_{t})\|^{2}]\leq\sigma^{2}.$
@@
(Non-convex convergence rate) Let Assumptions 1, 2, 3, 4, 5 hold and $q_{\omega}\mathrel{\mathop{:}}=(1+\omega)q<1$. Then, choosing $\eta=\frac{1}{\sqrt{T}}\leq\frac{\epsilon}{4LC_{0}}$, MicroAdam (Algorithm 4) satisfies
@@ Discussion.
First, notice that the leading term $\tfrac{1}{\sqrt{T}}$ of the rate is the optimal convergence speed for non-convex stochastic gradient methods (Ghadimi and Lan, 2016). Furthermore, the obtained convergence rate $\mathcal{O}(\tfrac{1}{\sqrt{T}}+\tfrac{d}{T})$ asymptotically matches the rate of uncompressed AMSGrad in the stochastic non-convex setup (Zhou et al., 2024a). Hence, the added compression framework of the MicroAdam together with error feedback mechanism can slow down the convergence speed only up to some constants. Evidently, the additional constants $C_{0}$ and $C_{2}$ affected by compression and appearing in the leading terms can be easily estimated once the compressors are fixed. Besides, if we store the full error information without applying $\mathcal{Q}$ compressor (i.e., $\omega=0,\,q_{\omega}=q$), then MicroAdam reduces to the single-node Comp-AMS method by Li et al. (2022) recovering the same convergence rate. The full proof is provided in the Appendix.
@@ 4.3 Convergence Rate for Non-Convex Functions under the PL Condition
(PL convergence rate) Let Assumptions 1, 2, 3, 4, 5 and 6 hold, and $q_{\omega}<1$. Then, choosing $\eta=\frac{C_{0}\log T}{\mu T}\leq\frac{\epsilon}{4LC_{0}}$, MicroAdam (Algorithm 4) satisfies
$\mathbb{E}[f(\theta_{T+1})]-f^{*}\leq\frac{f(\theta_{1})-f^{*}}{T}+\frac{\log T}{T}\left(\frac{LC_{0}^{2}}{\mu}\frac{\sigma^{2}+(C_{1}+C_{2}^{2})G^{2}}{\mu\epsilon}+\frac{C_{0}(1+C_{1})(1+d)G^{2}}{\mu\sqrt{\epsilon}}\right)\\ +\widetilde{\mathcal{O}}\left(\frac{G^{4}(G+d)}{T^{2}}\right)$
@@ 5 Experiments
We now validate our optimizer experimentally. We focus on comparing MicroAdam with Adam, Adam-8bit, GaLore and CAME in the context of language model finetuning on different tasks. Concretely, we test our optimizer in full finetuning (FFT) scenario on BERT-Base/Large (Devlin et al., 2018) and OPT-1.3B (Zhang et al., 2022) on GLUE/MNLI and Llama2-7B (Touvron et al., 2023) on the GSM8k math reasoning dataset and on the Open-Platypus instruction tuning dataset. We train all our models in bfloat16 format and tune the learning rates on a grid and report the best accuracy among 3 seeds. All methods use default AdamW parameters, such as $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$. MicroAdam uses a window size of $m=10$ gradients with $k=1\%$ density and quantization bucket size is set to 64 for the error feedback. For GaLore we use rank $r=256$ and the SVD update interval is set to $T=200$, as suggested by the original paper. We run our experiments on NVidia GPUs A100-SXM4-80GB and on RTX 3090 with 24GB RAM in single GPU setup. We provide full details regarding hyper-parameters in Appendix A.
Results on GLUE/MNLI.
We first test our integration of MicroAdam in HuggingFace Transformers (Wolf et al., 2020) on moderate-sized language models such as BERT-Base/Large (110M and 335M parameters) and OPT-1.3B, comparing with Adam, Adam-8bit, CAME and GaLore. The results are shown in Table 1. (Certain optimizers, notably CAME and GaLore, had numerical stability issues across runs; for a fair comparison, we report the numbers for the run with maximum accuracy. We emphasize that all methods were tuned using the same protocol.)
@@
From these results, we conclude that MicroAdam can provide better accuracy relative to other memory-efficient methods on moderate-sized models, at similar space costs. We show training loss curves in Appendix B.
@@
Results on LLaMA2-7B/GSM-8k.
Next, we perform finetuning on Llama-2 7B on GSM-8k, a challenging grade-school-level mathematical reasoning dataset. The baseline model obtains extremely low zero-shot accuracy on this task, therefore fine-tuning is necessary. In this setup, we compare MicroAdam with Adam and Adam-8bit in terms of evaluation accuracy and memory usage. In Table 2 we show our results for 3 training epochs, global batch size $32$ with micro-batch (per-device) size $1$, max sequence length $512$ on a single GPU. These parameters are standard for this task. We integrated our optimizer with the llm-foundry repository of MosaicML and tested via lm-evaluation-harness.
Table 2: FFT results for Llama-2 7B on GSM-8k
Adam Adam-8b MicroAdam
$m=10$ $m=20$
accuracy 34.50% 34.34% 34.72% 35.10%
memory 55.2 GB 42.5 GB 37.1 GB 39.7 GB
runtime 1h 17m 1h 18m 1h 8m 1h 37m
The results show that MicroAdam can allow accurate full fine-tuning of a 7B model on this task using a single 40GB GPU. Moreover, MicroAdam preserves accuracy relative to Adam, with lower memory usage than the well-optimized implementation of 8bit AdamW, and marginally lower running time for the shorter gradient window $m=10$. Increasing the window size $m$ to $20$ gradients leads to slightly better accuracy, at the cost of higher runtime and space. Running GaLore in this setup was infeasible since using SVD decomposition for all layers in the model was too slow; preliminary experiments (with high runtimes) did not yield competitive accuracy. We show training loss curves in Appendix B.
Results on LLaMA2-7B/Open-Platypus.
@@ Discussion.
In summary, the experimental results have shown that MicroAdam can recover the state-of-the-art accuracy of the the uncompressed Adam baseline, while providing significant memory gains and matching wall-clock speed on billion-parameter models. Specifically, our approach matches and outperforms Adam-8b and CAME both in terms of memory use and in terms of final accuracy. Relative to the high-compression GaLore method, MicroAdam provides consistently higher accuracy, as well as more stable practical convergence. We conclude that MicroAdam should be a good alternative to 8bit Adam in memory-constrained settings, and that the empirical results appear to validate our theoretical predictions.
@@ 6 Limitations and Broader Impact
The MicroAdam algorithm we propose is designed and tested with fine-tuning workloads in mind, where the user aims to minimize the memory cost of optimizing over a powerful pre-trained model. Additional work is needed to adapt our approach to the case of LLM pretraining, which presents a different set of challenges, both in terms of implementation and optimization trajectory. We plan to undertake this study in future work.
@@ Acknowledgements
Appendix A Hyper-parameters
In this section we provide details about the hyper-parameters that we used for each model and dataset. All Adam variants use default parameters $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8},\lambda=0$, where $\lambda$ is the weight decay parameter. We use $3$ seeds (7, 42 and 1234) for each learning rate and report the results for the best configuration that converged.
A.1 GLUE/MNLI
@@
A.2 GSM-8k.
For GSM-8k, we used the learning rate grid $\{1e-5,2e-5,3e-5,4e-5,5e-5,6e-7,7e-5,8e-5,9e-5\}$ and reported the model with the best evaluation accuracy. We found that different versions for PyTorch, lm-eval-harness and llm-foundry have large variance in the results.
@@
Appendix B Training Graphs
In this section we show training loss curves for BERT-Base, BERT-Large and OPT-1.3b on GLUE/MNLI and Llama-2 7B on GSM-8k.
@@ Appendix B Training Graphs
Appendix C Memory footprint for the optimizer state
@@
Appendix D Deferred Proofs
@@ D.1 Intermediate Lemmas
D.2 Non-convex Analysis
Here we derive the convergence rate with fixed step-size $\eta$. The rate shown in the main part can be obtained by plugging the expression of $\eta$. Note that the constants in the rate can be slightly improved if we choose $\eta=\min\{\frac{\epsilon}{4LC_{0}},\frac{1}{\sqrt{T}}\}$.
@@ D.2 Non-convex Analysis
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]\leq 2C_{0}\left(\frac{f(\theta_{1})-f^{*}}{T\eta}+\frac{\eta L\sigma^{2}}{\epsilon}+\frac{\eta LC_{2}^{2}G^{2}}{\epsilon}\right.\\ +\left.\frac{\eta^{2}LC_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}+\frac{(1+C_{1})G^{2}d}{T\sqrt{\epsilon}}+\frac{\eta(1+2C_{1})C_{1}LG^{2}d}{T\epsilon}\right),$
@@
$\overset{(b)}{\leq}\frac{\eta\rho}{2\epsilon}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{\eta^{3}L}{2\rho}\mathbb{E}\left[\left\|\frac{\frac{\beta_{1}}{1-\beta_{1}}m_{t-1}^{\prime}+\frac{\beta_{1}}{1-\beta_{1}}\mathcal{Z}_{t-1}+\mathcal{E}_{t}}{\sqrt{\hat{v}_{t-1}+\epsilon}}\right\|^{2}\right]+\frac{\eta^{2}C_{1}LG^{2}}{\sqrt{\epsilon}}\mathbb{E}[\|D_{t}\|_{1}]$
$\leq\frac{\eta\rho}{2\epsilon}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{\eta^{3}L}{2\rho}\frac{C_{1}^{2}G^{2}}{\epsilon}+\frac{\eta^{2}LC_{1}G^{2}}{\sqrt{\epsilon}}\mathbb{E}[\|D_{t}\|_{1}],$ (11)
@@
$IV$$\leq$ $\frac{\eta\rho}{2\epsilon}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{\eta^{3}L}{2\rho}\frac{C_{1}^{2}G^{2}}{\epsilon}+\frac{\eta^{2}LC_{1}G^{2}}{\sqrt{\epsilon}}\mathbb{E}[\|D_{t}\|_{1}],$
@@
$\leq\left(-\frac{\eta}{C_{0}}+\frac{\eta^{2}L}{\epsilon}+\frac{\eta\rho}{2\epsilon}\right)\sum_{t=1}^{T}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{T\eta^{2}L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon}+\frac{T\eta^{3}LC_{1}^{2}G^{2}}{2\rho\epsilon}$
@@
$\quad+\frac{T\eta^{3}LC_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}+\frac{\eta(1+C_{1})G^{2}d}{\sqrt{\epsilon}}+\frac{\eta^{2}(1+2C_{1})C_{1}LG^{2}d}{\epsilon}.$
@@
$\leq 2C_{0}\left(\frac{\mathbb{E}[f(x_{1})-f(x_{T+1})]}{T\eta}+\frac{\eta L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon}+\frac{\eta^{2}LC_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}\right)$
@@
$\leq 2C_{0}\left(\frac{f(\theta_{1})-f^{*}}{T\eta}+\frac{\eta L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon}+\frac{\eta^{2}LC_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}\right)$
@@
D.3 Analysis Under PL Condition
As in the non-convex analysis, here we derive the convergence rate with fixed step-size $\eta$. The rate shown in the main part can be obtained by plugging the expression of $\eta$. Also, the constants in the rate can be slightly improved if we choose $\eta=\min\{\frac{\epsilon}{4LC_{0}},\frac{C_{0}\log T}{\mu T}\}$.

### v2

@@ 4.1 Gradient and Error Compression
9:     $\theta_{t+1}=\theta_{t}-\eta\frac{{\color[rgb]{0,0,0}m_{t}}}{\sqrt{\hat{v}_{t}+\epsilon}}$diamond Update the model parameters
@@ 4.2 Convergence Guarantees for General Smooth Non-convex Functions
The loss function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is lower bounded by some $f^{*}\in\mathbb{R}$ and $L$-smooth, i.e., $\left\|\nabla f(\theta)-\nabla f(\theta^{\prime})\right\|\leq L\left\|\theta-\theta^{\prime}\right\|$, for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}.$
@@
For all iterates $t\geq 1$, the stochastic gradient $g_{t}$ is unbiased and uniformly bounded by a constant $G\geq 0$, i.e., $\mathbb{E}[g_{t}]=\nabla f(\theta_{t}),\;\left\|g_{t}\right\|\leq G.$
@@
For all iterates $t\geq 1$, the variance of the stochastic gradient $g_{t}$ is uniformly bounded by some constant $\sigma^{2}\geq 0$, i.e., $\mathbb{E}[\|g_{t}-\nabla f(\theta_{t})\|^{2}]\leq\sigma^{2}.$
@@
(Non-convex convergence rate) Let Assumptions 1, 2, 3, 4, 5 hold and $q_{\omega}\mathrel{\mathop{:}}=(1+\omega)q<1$. Then, choosing ${\color[rgb]{0,0,0}\eta=\min\{\frac{\epsilon}{4LC_{0}},\frac{1}{\sqrt{T}}\}}$, MicroAdam (Algorithm 3) satisfies
@@ Discussion.
First, notice that the leading term $\tfrac{1}{\sqrt{T}}$ of the rate is the optimal convergence speed for non-convex stochastic gradient methods (Ghadimi and Lan, 2016). Furthermore, the obtained convergence rate $\mathcal{O}(\tfrac{1}{\sqrt{T}}+\tfrac{d}{T})$ asymptotically matches the rate of uncompressed AMSGrad in the stochastic non-convex setup (Zhou et al., 2024a). Hence, the added compression framework of the MicroAdam together with error feedback mechanism can slow down the convergence speed only up to some constants including the dimension. Evidently, the additional constants $C_{0}$ and $C_{2}$ affected by compression and appearing in the leading terms can be easily estimated once the compressors are fixed. Besides, if we store the full error information without applying $\mathcal{Q}$ compressor (i.e., $\omega=0,\,q_{\omega}=q$), then MicroAdam reduces to the single-node Comp-AMS method by Li et al. (2022) recovering the same convergence rate. The full proof is provided in the Appendix.
@@ 4.3 Convergence Rate for Non-Convex Functions under the PL Condition
(PL convergence rate) Let Assumptions 1, 2, 3, 4, 5 and 6 hold, and $q_{\omega}<1$. Then, choosing ${\color[rgb]{0,0,0}\eta=\min\{\frac{\epsilon}{4LC_{0}},\frac{2C_{0}\log T}{\mu T}\}}$, MicroAdam (Algorithm 3) satisfies
$\textstyle\mathbb{E}[f(\theta_{T+1})]-f^{*}\leq\frac{{\color[rgb]{0,0,0}2}\log T}{T}\left(\frac{LC_{0}^{2}}{\mu}\frac{\sigma^{2}+(C_{1}+C_{2}^{2})G^{2}}{\mu\epsilon}+\frac{C_{0}(1+C_{1})(1+d)G^{2}}{\mu\sqrt{\epsilon}}\right)+\widetilde{\mathcal{O}}\left(\frac{G^{4}(G+d)}{T^{2}}\right)$
@@ 5 Experiments
We now validate our optimizer experimentally. We focus on comparing MicroAdam with Adam, Adam-8bit, GaLore and CAME in the context of LLM finetuning on different tasks and with SGD, Adam and AdamW-8bit in the context of ResNets on ImageNet. Concretely, we test our optimizer in full finetuning (FFT) scenario on BERT-Base/Large (Devlin et al., 2018) and OPT-1.3B (Zhang et al., 2022) on GLUE/MNLI and Llama2-7B/13B (Touvron et al., 2023) on the GSM8k math reasoning dataset and on the Open-Platypus instruction tuning dataset, as well as pre-training ResNet models on ImageNet. We provide full details regarding training settings hyper-parameters in Appendix B.
Finetuning results on GLUE/MNLI.
We first test our integration of MicroAdam in HuggingFace Transformers (Wolf et al., 2020) on moderate-sized language models such as BERT-Base/Large (110M and 335M parameters) and OPT-1.3B, comparing with Adam, Adam-8bit, CAME and GaLore. The results are shown in Table 1. Certain optimizers, notably CAME and GaLore, had numerical stability issues across runs; for a fair comparison, we report the numbers for the run with maximum accuracy. We emphasize that all methods were tuned using the same protocol.
@@
From these results, we conclude that MicroAdam can provide better accuracy relative to other memory-efficient methods on moderate-sized models, at similar space costs. We show training loss curves in Appendix C.
@@
Finetuning results for LLaMA2 on GSM-8k.
Next, we perform finetuning on Llama-2 7B/13B on GSM-8k, a challenging grade-school-level mathematical reasoning dataset. The baseline model obtains extremely low zero-shot accuracy on this task and therefore fine-tuning is necessary. In this setup, we compare MicroAdam with Adam and Adam-8bit in terms of evaluation accuracy and memory usage. In Table 2 we show our results for 3 training epochs, global batch size $32$ with micro-batch (per-device) size $1$, max sequence length $512$ on a single GPU, which are the standard parameters for this task. We integrated our optimizer with the llm-foundry repository of MosaicML and tested via lm-evaluation-harness.
For the 7B model, out results show that MicroAdam can allow accurate full fine-tuning of a 7B model on this task using a single 40GB GPU. Moreover, MicroAdam preserves accuracy relative to Adam, with lower memory usage than the well-optimized implementation of 8bit AdamW, and marginally lower running time for the shorter gradient window $m=10$. Increasing the window size $m$ to $20$ gradients leads to slightly better accuracy, at the cost of higher runtime and space, but still in the 40GB limit. Running GaLore in this setup was infeasible since using SVD decomposition for all layers in the model was too slow. Preliminary experiments (with high runtimes) did not yield competitive accuracy. We show training loss curves in Appendix C.
Table 2: FFT results for Llama-2 7B/13B on GSM-8k.
LLaMA-2 size Optimizer Accuracy State Total Runtime
Adam 34.50% 25.1 GB 55.2 GB 1h 17m
7B Adam-8b 34.34% 12.55 GB 42.5 GB 1h 18m
MicroAdam ($m=10$) 34.72% 5.65 GB 37.1 GB 1h 8m
MicroAdam ($m=20$) 35.10% 8.25 GB 39.7 GB 1h 37m
Adam 47.08% 48.42 GB >80 GB 1h 20m
13B Adam-8b 45.19% 24.21 GB >80 GB 1h 17m
MicroAdam ($m=10$) 44.88 % 10.9 GB 70 GB 1h 38m
Finetuning results for LLaMA2-7B on Open-Platypus.
@@
Pre-training results for ResNets on ImageNet.
In Table 4 we present our results for pre-training (from scratch, randomly initialized weights) for ResNet-18/50 on ImageNet (see Figure 6 and Figure 7 in Appendix C). We compare our MicroAdam with SGD, Adam and AdamW and report the training loss, validation accuracy and only the optimizer state memory because the total memory usage is not an issue for ResNets on ImageNet (here we focus on the results to emphasize the pre-training performance). For ResNet-18, AdamW reaches the lowest training loss, followed by AdamW-8bit, MicroAdam and in the end MicroAdam. However, despite slightly larger loss, MicroAdam yields the best result among all optimizers, with  2% more than the highly tuned SGD, while having the lowest memory footprint for the optimizer states. For ResNet-50, AdamW-8bit reaches the lowest training loss, followed by AdamW, MicroAdam and SGD. The validation accuracy for AdamW and AdamW-8bit is surprisingly small compared to SGD and MicroAdam. As it is widely known in the community, Adam variants have lower performance than SGD for Computer Vision tasks and MicroAdam fixes this issue (see the Discussion section for an intuitive explanation for this phenomenon).
Table 4: Pre-training results for ResNet-18 and ResNet-50 on ImageNet.
Model Metric SGD AdamW AdamW-8bit MicroAdam
Train Loss 1.416 1.087 1.104 1.218
ResNet-18 Accuracy 69.96% 69.83% 70.13% 71.86%
State Size 44.59 MB 89.18 MB 22.30 MB 10.03 MB
Train Loss 0.9770 0.5344 0.5158 0.7732
ResNet-50 Accuracy 76.24% 72.05% 72.48% 77.37%
State Size 97.49 MB 194.98 MB 48.75 MB 21.94 MB
Pre-training LLMs.
In this section we explain why we do not include LLM pre-training results. Our motivation is twofold. First, MicroAdam is mainly designed for low-memory finetuning, and the experimental section shows that MicroAdam achieved this goal. Surprisingly, updating only 10% of the weights at each step yields to significantly better performance compared to SGD for ResNets on ImageNet. Secondly, our experiments on LLM pre-training showed difficulties in achieving the same performance compared to AdamW-8bit. Our explanation is that projection matrices from the attention layers must receive dense updates to learn the correlations between words. In contrast to the convolutional filters for CV models, the weights in attention are much larger and try to capture global correlations (features) between words, while the convolutional filters are smaller and capture local features.
@@ Discussion.
In Appendix A we provide information about the optimization set, intuitive explanations for the implicit regularization effect of MicroAdam, as well as an overview of our results.
@@ 6 Limitations and Broader Impact
The MicroAdam algorithm we propose is designed and tested with fine-tuning workloads in mind, where the user aims to minimize the memory cost of optimizing over a powerful pre-trained model. Additional work is needed to adapt our approach to the case of LLM pre-training, which presents a different set of challenges, both in terms of implementation and optimization trajectory. We plan to undertake this study in future work as the current implementation works for ResNets.
@@ Acknowledgements
Appendix A Additional Explanations and Experimental Details
Optimization set.
The parameters included in the optimization set usually vary depending on the model and optimizer type. For GLUE, we do not include the embeddings in the optimization set for any of the optimizers because our experiments showed no significant difference when optimizing the embeddings. Moreover, it is more fair for GaLore which would have an increased memory usage due to the projection matrix for the embeddings. For LLaMa2 and ResNet models, we include all layers in the optimization set, regardless of their type. This means that MicroAdam updates at most $10\%$ of the weights in each layer. In the original work, GaLore was applied only to a subset of layers, such as Q-, K-, V-, O-, up-, down- and gate-projection layers. Applying GaLore to all layers in the same way as we do with MicroAdam would result in much larger memory usage because of the projection matrices for the embeddings. Moreover, computing SVD for the embedding layer would be infeasible. As a result, we omit the GaLore for LLaMa2 and ResNet experiments.
Implicit regularization.
In our results so far, we observed MicroAdam having better performance for a few LLM finetuning tasks, but especially for the ResNet/ImageNet results, where the difference was statistically significant (around 1%). Our intuition for this behavior mainly comes from the accuracy curves in Figure 6 and Figure 7. The trajectories of MicroAdam and SGD are typical for regularized training. We hypothesize that MicroAdam has an implicit regularization mechanism because the model update $u_{t}$ is  90% sparse, which leads to only updating  10% of the model parameters at each step in each layer. In contrast to a 100% dense update $u_{t}$, a sparse update would not change the model parameters as much as a dense one. In the ResNet experiments, all optimizers used the same regularization parameter $\lambda=1e-4$, but the accuracy graph shows that MicroAdam is more regularized than SGD, while the graphs for AdamW and AdamW-8bit look like the regularization does not have any effect.
Discussion.
In summary, the experimental results have shown that MicroAdam can recover the state-of-the-art accuracy of the the uncompressed Adam baseline, while providing significant memory gains and matching wall-clock speed on billion-parameter models. Specifically, our approach matches and outperforms Adam-8b and CAME both in terms of memory use and in terms of final accuracy. Relative to the high-compression GaLore method, MicroAdam provides consistently higher accuracy, as well as more stable practical convergence. We conclude that MicroAdam should be a good alternative to Adam-8bit in memory-constrained settings, and that the empirical results appear to validate our theoretical predictions.
Appendix B Training Settings and Hyper-parameters
In this section we provide details about the hyper-parameters that we used for each model and dataset. We train all our models in bfloat16 format, tune the learning rates on a grid and report the best accuracy among 3 seeds (7, 42 and 1234) and report the results for the best configuration that converged.
All Adam variants use default parameters $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$ and the regularization parameter $\lambda$ is 00 for finetuning and $43e-$ 4 for ImageNet pre-training. MicroAdam uses a window size of $m=10$ gradients with $k=1\%$ density (equivalent to $99\%$ sparsity and quantization bucket size is set to 64 for the error feedback.
For GaLore we use rank $r=256$ and the SVD update interval is set to $T=200$, as suggested by the original paper. We run our experiments on NVidia GPUs A100-SXM4-80GB, H100-80GB and on RTX 3090 with 24GB RAM in single GPU setup.
B.1 GLUE/MNLI
@@
B.2 GSM-8k.
For GSM-8k, we used the learning rate grid $\{1e-5,2e-5,3e-5,4e-5,5e-5,6e-5,7e-5,8e-5,9e-5\}$ and reported the model with the best evaluation accuracy. We found that different versions for PyTorch, lm-eval-harness and llm-foundry have large variance in the results.
@@
B.3 ImageNet
For ImageNet, we integrate our MicroAdam in the FFCV repository, which is highly tuned for ResNets and SGD. We use $E=100$ epochs, batch size 1024, cosine learning rate schedule with warmup and image resolution $224\times 224$ and precision bfloat16. We started from the initial learning rate $\eta=1.024$ tuned in the repository which scored highest accuracy for SGD. This learning rate also worked well for MicroAdam, but it didn’t work for AdamW and AdamW-8bit. For these two Adam variants we divided the learning rate by 2 until the models converged.
ResNet-18.
For AdamW, the learning rate is $\eta=0.016$ and for AdamW-8bit is $\eta=0.032$.
ResNet-50.
For AdamW, the learning rate is $\eta=0.008$ and for AdamW-8bit is $\eta=0.008$.
Appendix C Training Graphs
In this section we show training loss curves for BERT-Base, BERT-Large and OPT-1.3b on GLUE/MNLI and Llama-2 7B/13B on GSM-8k and ResNet-18/50 on ImageNet.
@@ Appendix C Training Graphs
Figure 6: Pre-training for ResNet-18 on ImageNet
Figure 7: Pre-training for ResNet-50 on ImageNet
Appendix D Memory footprint for the optimizer state
@@
Appendix E Deferred Proofs
@@ E.1 Intermediate Lemmas
E.2 Non-convex Analysis
Here we derive the convergence rate with fixed step-size $\eta$. The rate shown in the main part can be obtained by plugging the expression of $\eta$ shown after the proof.
@@ E.2 Non-convex Analysis
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]\leq 2C_{0}\left(\frac{f(\theta_{1})-f^{*}}{T\eta}+\frac{\eta L\sigma^{2}}{\epsilon}+\frac{\eta LC_{2}^{2}G^{2}}{\epsilon}\right.\\ \textstyle+\left.\frac{\eta^{2}L^{{\color[rgb]{0,0,0}2}}C_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}+\frac{(1+C_{1})G^{2}d}{T\sqrt{\epsilon}}+\frac{\eta(1+2C_{1})C_{1}LG^{2}d}{T\epsilon}\right),$
@@
$\overset{(b)}{\leq}\frac{\eta\rho}{2\epsilon}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{\eta^{3}L^{{\color[rgb]{0,0,0}2}}}{2\rho}\mathbb{E}\left[\left\|\frac{\frac{\beta_{1}}{1-\beta_{1}}m_{t-1}^{\prime}+\frac{\beta_{1}}{1-\beta_{1}}\mathcal{Z}_{t-1}+\mathcal{E}_{t}}{\sqrt{\hat{v}_{t-1}+\epsilon}}\right\|^{2}\right]+\frac{\eta^{2}C_{1}LG^{2}}{\sqrt{\epsilon}}\mathbb{E}[\|D_{t}\|_{1}]$
$\leq\frac{\eta\rho}{2\epsilon}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{\eta^{3}L^{{\color[rgb]{0,0,0}2}}}{2\rho}\frac{C_{1}^{2}G^{2}}{\epsilon}+\frac{\eta^{2}LC_{1}G^{2}}{\sqrt{\epsilon}}\mathbb{E}[\|D_{t}\|_{1}],$ (11)
@@
$IV$$\leq$ $\frac{\eta\rho}{2\epsilon}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{\eta^{3}L^{{\color[rgb]{0,0,0}2}}}{2\rho}\frac{C_{1}^{2}G^{2}}{\epsilon}+\frac{\eta^{2}LC_{1}G^{2}}{\sqrt{\epsilon}}\mathbb{E}[\|D_{t}\|_{1}],$
@@
$\leq\left(-\frac{\eta}{C_{0}}+\frac{\eta^{2}L}{\epsilon}+\frac{\eta\rho}{2\epsilon}\right)\sum_{t=1}^{T}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]+\frac{T\eta^{2}L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon}+\frac{T\eta^{3}L^{{\color[rgb]{0,0,0}2}}C_{1}^{2}G^{2}}{2\rho\epsilon}$
@@
$\quad+\frac{T\eta^{3}L^{{\color[rgb]{0,0,0}2}}C_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}+\frac{\eta(1+C_{1})G^{2}d}{\sqrt{\epsilon}}+\frac{\eta^{2}(1+2C_{1})C_{1}LG^{2}d}{\epsilon}.$
@@
$\leq 2C_{0}\left(\frac{\mathbb{E}[f(x_{1})-f(x_{T+1})]}{T\eta}+\frac{\eta L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon}+\frac{\eta^{2}L^{{\color[rgb]{0,0,0}2}}C_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}\right)$
@@
$\leq 2C_{0}\left(\frac{f(\theta_{1})-f^{*}}{T\eta}+\frac{\eta L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon}+\frac{\eta^{2}L^{{\color[rgb]{0,0,0}2}}C_{0}C_{1}^{2}G^{2}}{\epsilon^{2}}\right)$
@@
To get the rate mentioned in the main part, choose $\eta=\min\{\frac{\epsilon}{4LC_{0}},\frac{1}{\sqrt{T}}\}$ and continue
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\|\nabla f(\theta_{t})\|^{2}]$
$\leq 2C_{0}\left(\max\left\{1,\frac{4LC_{0}}{\epsilon\sqrt{T}}\right\}\frac{f(\theta_{1})-f^{*}}{\sqrt{T}}+\frac{L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon\sqrt{T}}+\frac{L^{2}C_{0}C_{1}^{2}G^{2}}{\epsilon^{2}T}\right)$
$\quad+2C_{0}\left(\frac{(1+C_{1})G^{2}d}{T\sqrt{\epsilon}}+\frac{(1+2C_{1})C_{1}LG^{2}d}{\epsilon T^{3/2}}\right)$
$\leq 2C_{0}\left(\frac{f(\theta_{1})-f^{*}}{\sqrt{T}}+\frac{L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon\sqrt{T}}\right)$
$\quad+2C_{0}\left(\frac{4LC_{0}}{\epsilon}\frac{f(\theta_{1})-f^{*}}{T}+\frac{L^{2}C_{0}C_{1}^{2}G^{2}}{\epsilon^{2}T}+\frac{(1+C_{1})G^{2}d}{T\sqrt{\epsilon}}+\frac{(1+2C_{1})C_{1}LG^{2}d}{\epsilon T^{3/2}}\right)$
$=2C_{0}\left(\frac{f(\theta_{1})-f^{*}}{\sqrt{T}}+\frac{L(\sigma^{2}+C_{2}^{2}G^{2})}{\epsilon\sqrt{T}}\right)+\mathcal{O}\left(\frac{G^{3}(G+d)}{T}\right),$
where in the second part of the rate we suppressed all the problem and compression dependent constants.
E.3 Analysis Under PL Condition
As in the non-convex analysis, here we derive the convergence rate with fixed step-size $\eta$. The rate shown in the main part can be obtained by plugging the expression of $\eta$.